40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5111     Entry time: Wed Aug 3 16:18:30 2011
Author: steve 
Type: Update 
Category: VAC 
Subject: vent is almost finished 

P1 at 750 Torr,  Vent valve VV1 is left open to atm. It will reach equilibrium in an hour. The PSL power is not turned down yet. The beam path is blocked at two places inside the enclosure.

The low power 20 mW  input  MC can be locked tonight.

 

Next day's - Thursday notes:

We  reached 760 Torr of atm in 8 hrs at  1.7 Torr/ min with 5 cylinders of  instrument grade air.

Used 500 PSI of nitrogen to get to 20 Torr at first.  This flow  and the first cylinder of N2  has to be slower next time.

 

We did the following:

1, oplevs were centered on locked caveties 2 days ago

2, just before vent we rechecked arms  pointing, they wre still locking

3,eloged strain gages  and turned HVs off at PZT-steering   at 1Y3 rack and OMC-PZT ps at 1X2 aux rack. They were actually off.

4,,checked jam nuts 6 places x 3

5, turned oplev servos off

6, eloged SUS summery page

7, checked particle counts at ITMY: 1 micron 300,  0.5 micron 3K,  0.3 micron  30K particles / cft-min. All three cranes were wiped 2 days ago.

8, blocked laser into IFO: main 1064 in 1.25W were blocked by manual 2 blocks on PSL table. The PSL output shutter was left open to be triggered by interlock  at P1 3 mTorr. It did!

    Aux NPRO-AP and green light  input at both ends were blocked.

Attachment 1: vent8h.jpg  35 kB  Uploaded Thu Aug 4 16:54:44 2011  | Hide | Hide all
vent8h.jpg
ELOG V3.1.3-